# DynamicPAE: Generating Scene-Aware Physical Adversarial Examples in Real-Time

### v1

@@ 5 Related Works
Physical Adversarial Attacks: Based on the paradigm of optimizing static PAEs, research focuses on simulating the physical world [69] and the target AI system [20], and modeling PAEs under naturalness constrains [42] or new attack mediums [70]. Recent works began to focus on the challenge of generating dynamic PAEs. [24] proposes the Dynamic Adversarial Patch on the dynamically changing clothes, while the PAE data itself is still static. [23] proposes the adaptive PAE by manually clustering the attack scenario and optimizing the static PAE for each cluster, but it is not suitable for the open world. [22] proposes a physical adversarial attack by controlling a dynamic laser beam in the simulation using reinforcement learning. Nevertheless, it is only capable of modeling a limited number of laser states but not the general space of PAE.
@@
Generative Models and Generative Adversarial Attacks: Motivated by learning representations [71], classical generative neural networks adapt generative learning methods to train a model that maps the space of real data $\mathcal{X}$ to a latent space $\mathcal{Z}$, which has better mapping modeling and sampling efficiency. With refined learning tasks and neural network construction [49, 72], generative models is able to generate complex data, e.g., model the relation of texts and natural images [73], molecular dynamics [74], protein structure [75] and images and corresponding digital perturbation AEs [37]. In terms of adversarial attacks, NN-based generators have been applied in improving perturbation optimization [37], generating naturalistic AEs by modeling the data manifold [68, 21, 42, 76] and probabilistic modeling perturbations for black-box attacks [77]. However, only unconditional & static PAEs have been proposed under the generative framework [42, 78, 76]. Except for some specific adversarial attack mediums, e.g., audio [79] and face subspace [70], the physical-digital gap [18] is small, the randomized transformation may not be necessary, and therefore the digital AE generation network may be effective. As an important step forward, we bridge the general gap between scene-aware dynamic PAE and generative NNs.
@@
Multidisciplinary Optimization Techniques: Similar optimization problems also occurred in the application of neural networks [80] that learns complex tasks in the field of AI4Science, e.g., approximating complex fields [81] and training on noisy measurements [82]. We believe the relevant research will be beneficial for scene-aware PAE generation, and visa versa. In the field of reinforcement learning, research has been conducted on the problem of exploring the action space and utilizing inaccurate feedback [83]. However, the general scene-aware PAE generation problem is defined in the complex state space, such as a patch with 10K+ pixels, which cannot is intractable for reinforcement models. Recently, language-driven models have been adapted for optimization [84], but there is still a gap between language and low-level AE data [85], which should be filled by the generative embedding model.
6 Conclusions
@@
Future research directions can focus on technical improvements to the patch representation and training mechanism, exploring the application of the proposed framework, and facilitating the developed techniques in broader domains. In terms of application, although we have presented the preliminary prototype in our experiments, work on generating adversarial test and training data and exploring real-world red team attacks for specific tasks is still worthy of further investigation. We believe that our study has the broader potential to benefit the optimization of defense-insensitive adversarial attacks and noisy open-world tasks, and the model for the hardness of PAE optimization can bring insights to the defense of PAEs. In addition, the proposed scene-aware PAE generation task is representative and beneficial for foundation model research.
@@ Acknowledgments
This work is supported by Zhongguancun Laboratory and State Key Laboratory of Complex & Critical Software Environment (CCSE).

### v3

@@ 4.5 Ablation Studies
TABLE XVII: The ablation of skewness alignment. $\lambda_{\textrm{Atk}}$ is set by the deployment scenario. By controlling the skewness $\varsigma$ of the stealth distribution of generated PAEs in training, better extreme attack performance and custom stealthiness configuration can be achieved during deployment. Stealthiness is evaluated using SSIM (higher is better).
$\varsigma_{\text{target}}=-1$ $\varsigma_{\text{target}}=+1$ $\varsigma_{\text{target}}=0$
$\lambda_{\text{Atk}}$ $AP_{50}$ SSIM↑ $AP_{50}$ SSIM↑ $AP_{50}$ SSIM↑
Yolo-v5s
0.0 77.9% 0.99385 77.8% 0.99394 78.2% 0.99387
0.2 75.9% 0.99380 50.7% 0.99326 74.4% 0.99385
0.5 50.3% 0.99329 24.2% 0.99206 43.8% 0.99304
0.8 25.7% 0.99205 16.5% 0.99031 22.5% 0.99172
1.0 16.8% 0.98913 14.9% 0.98904 14.9% 0.98905
Yolo-v8n
0.0 85.7% 0.99385 85.8% 0.99398 85.7% 0.99375
0.2 78.0% 0.99374 77.9% 0.99353 79.5% 0.99371
0.5 68.4% 0.99347 42.6% 0.99169 58.8% 0.99293
0.8 34.4% 0.99120 34.0% 0.99048 33.1% 0.99118
1.0 29.4% 0.98918 32.9% 0.98949 26.0% 0.98925
$\bullet$ Latent Regularization: Due to the unique characteristic of DynamicPAE Training, additional regularization is applied on the latent encoding to prevent gradient explosion. We compare the differences in regularization in terms of the attack performance. As shown in Table XVI, the center-regularized $\boldsymbol{z}_{adv}$ achieves the optimal performance among different settings and solves the problem of gradient explosion when learning diverse PAEs through the auxiliary residual task. Note that the PAE decoder $\mathrm{Dec}$ degenerates to a single static PAE without the auxiliary residual task. We noticed classical VAE regularization still suffers the infinite gradient in the model of Yolo-v8n and the attack performance on Yolo-v5 is lower, indicating the widely-used Gaussian prior might not be a proper selection for the latent of explored PAEs. We further analyze the gradient explosion problem in supplementary materials.
$\bullet$ Skewness Alignment: The proposed skewness-aligned objective re-weighting controls the stealthiness distribution of PAEs when different $\lambda_{\mathrm{Atk}}$ is set, and it is also beneficial for improving the attack performance. We perform the ablation study by altering the target skewness in Eq 15. The results are shown in Table XVII, indicating that (1) neutral setting of target skewness ($\varsigma_{\text{target}}=0$) achieves the highest extreme attack performance when $\lambda_{\mathrm{Atk}}=1$. This shows that balancing the attack and residual task is beneficial for exploring more aggressive PAE patterns, and the precise balance indicator is worth further study. (2) After deployment, the user can adjust the stealthiness by setting different $\lambda_{\mathrm{Atk}}$ as the input of the generator $\mathcal{G}$, and the generator can select the proper PAE pattern based on $\lambda_{\mathrm{Atk}}$
4.6 Applications and Defenses
This section discusses the use of the proposed attack method in adversarial test data generation, along with potential defense strategies.
$\bullet$ Application on Simulated Testing: To evaluate the applicability of the proposed model in the simulated testing, we obtain the simulation results from different perspectives in the CARLA autonomous driving simulation platform [62], divide them into training and testing datasets, and apply the adversarial example as a perturbation of the vehicle’s texture. Supplement to the digital environment benchmark, we compare the performance with the GAN-based perturbation generation framework AdvGAN, since both our model and AdvGAN are learning-based and real-time. The adversarial attack scenario is shown in Figure 15. In the Identical View scenario, the attacker’s observation, as the generative attacker’s input, is identical to the victim model. In the Transformed View scenario, the view of the victim model is on the other side. Following the Conditional-Uncertainty-Aligned Data Model, we formulate the distribution $\Omega$ in Eq. 12 that generates $\mathbf{P}_{\mathrm{X}}$ as:
[…]
where $\mathcal{T}$ is the view transformation performed by the simulation engine, $\mathrm{Cam}_{X}$ denotes the sampled camera state.
The results are in Table XVIII, showing that besides patch attacks, our model could also achieve strong perturbation attack performance, outperforming AdvGAN. To measure the stealthiness of the attack more accurately, we measure the SSIM similarity in addition to the $l_{\infty}$ constraint. Interestingly, our method performs better in the transformed view than in the identical view when $\varepsilon=1\%$, which may be attributed to the increased generalization due to the augmented transformation $\mathcal{T}$. This suggests our method is more suitable for complex physical attacks. We further test the robustness of the perturbation under extreme weather conditions by adding a snow & rain simulation provided by the Kornia library [63] during the test. The results show that (1) only adding extreme weather leads to a mild decrease in detection precision. (2) DynamicPAE maintains significant attack performance even after the perturbation is transformed by the simulation, and the performance may even increase due to the collaboration of multiple distortions. These results demonstrate DynamicPAE’s cross-environment robustness, indicating its strong extensibility.
TABLE XVIII: Results on CARLA simulation. The illustration of different view settings is shown in Figure 15. Ex indicates integrating extreme weather simulation.
Identical View Transformed View
$\varepsilon$ Method AP50 AP01 SSIM AP50 AP01 SSIM
0% Clean 43.23% 46.56% 1.0000 51.08% 52.56% 1.0000
Ex. weather 35.35% 43.88% 0.7833 44.35% 50.35% 0.7475
1% AdvGAN 29.67% 38.40% 0.9983 34.30% 44.17% 0.9994
1% DynamicPAE 5.35% 12.39% 0.9989 3.31% 7.53% 0.9994
1% DynamicPAE-Ex 1.56% 10.43% 0.7823 3.95% 8.11% 0.7466
5% AdvGAN 6.0e-5 7.0e-3 0.9888 2.9e-3 1.1e-2 0.9928
5% DynamicPAE <<1e-6 2.5e-6 0.9969 <<1e-6 8.5e-6 0.9931
5% DynamicPAE-Ex <<1e-6 <<1e-6 0.7752 <<1e-6 <<1e-6 0.7418
$\bullet$ Countermeasure on the Attack: Since this study aims to generate PAEs, which might threaten real-world deployed AI applications, it is necessary to provide possible countermeasures to make it responsible. We provide the experimental results by a typical adversarial defense approach, i.e., adversarial patch detection. Note that we believe that the adversarial training methods could also serve as a useful defense against dynamic adversarial patch attacks in principle, but we leave it as an open problem.
Specifically, the adversarial patch detection methods try to train an object detector to locate the dangerous patches inside AEs. Therefore, we construct a test dataset based on COCO-val with PAEs injected. We then introduce a SOTA method, namely, NAPGuard [57], as a defender with 2 different evaluation settings, i.e., the patch generated by attacking Yolo-v5s and Yolo-v8n. We select the Ultralytics Yolov5s as the backbone for NAPGuard. For our framework, we provide 3 parameter settings for the evaluation, to wit, setting $\lambda=0.5,0.6,1.0$ (DynamicPAEA in Table V adopts $\lambda=1.0$). As for metrics, we report the AP50 of the object detector and the adversarial patch detector (denoted as “Attack AP50” & “Defense AP50”), respectively. The evaluation results can be witnessed in Table XIX, where we can observe the following conclusions: ❶ compared with the GAN-NAP and T-sea, the NAPGuard could effectively detect DynamicPAEλ=1.0, i.e., its Defense AP50 achieves 0.9534, while that of the GAN-NAP and T-sea is respectively 0.6597 and 0.7099. That means the most aggressive settings($\lambda=1.0$) might be detected more easily, which shows that its greatest potential social risk is in control. ❷ The tendency of the evaluation results in different settings shows that our proposed method could sacrifice aggressiveness for stealthiness, indicating its dynamic attacking ability in the physical world. Therefore, it is worth studying targeted defense approaches against this typical attack. ❸ For different target models, i.e., Yolo-v5s and Yolo-v8n, our framework shows defense-evading ability, which may be attributed to the generalizability of the defense method, urging the study of more effective and efficient countermeasures.
TABLE XIX: Performance of countermeasure methods.
Attack AP50 Defense AP50
Method Yolo-v5s Yolo-v8n Yolo-v5s Yolo-v8n
GAN-NAP 0.4014 0.7103 0.6597 0.9769
T-sea 0.4825 0.6043 0.7099 0.9787
DynamicPAEλ=0.5 0.4375 0.5875 0.6709 0.2024
DynamicPAEλ=0.6 0.3630 0.4653 0.7537 0.3232
DynamicPAEλ=1.0 0.1487 0.2604 0.8307 0.6609
@@ 5 Related Works
Physical Adversarial Attacks: Based on the paradigm of optimizing static PAEs, research focuses on simulating the physical world [68] and the target AI system [13], and modeling PAEs under naturalness constrains [37] or new attack mediums [69]. Recent works began to focus on the challenge of generating dynamic PAEs. [17] proposes the Dynamic Adversarial Patch on the dynamically changing clothes, while the PAE data itself is still static. [16] proposes the adaptive PAE by manually clustering the attack scenario and optimizing the static PAE for each cluster, but it is not suitable for the open world. [15] proposes a physical adversarial attack by controlling a dynamic laser beam in the simulation using reinforcement learning. Nevertheless, it is only capable of modeling a limited number of laser states but not the general space of PAE. [70] proposes a computational efficient optimization and rendering pipeline for 3D physical attacks. However, the dynamic adaptation to the current scene is left as an open problem.
@@
Generative Models and Generative Adversarial Attacks: Motivated by learning representations [71], classical generative neural networks adapt generative learning methods to train a model that maps the space of real data $\mathcal{X}$ to a latent space $\mathcal{Z}$, which has better mapping modeling and sampling efficiency. With refined learning tasks and neural network construction [41, 42], generative models are able to generate complex data, e.g., model the relation of texts and natural images [72], molecular dynamics [73], protein structure [74] and images and corresponding digital perturbation AEs [31]. In terms of adversarial attacks, NN-based generators have been applied in improving perturbation optimization [31], generating naturalistic AEs by modeling the data manifold [67, 14, 37, 75] and probabilistic modeling perturbations for black-box attacks [76]. However, only unconditional & static PAEs have been proposed under the generative framework [37, 77, 75]. Except for some specific adversarial attack mediums, e.g., audio [78] and face subspace [69], the physical-digital gap [22] is small, the randomized transformation may not be necessary, and therefore the digital AE generation network may be effective. As an important step forward, we bridge the general gap between scene-aware dynamic PAE and generative NNs.
@@
Multidisciplinary Optimization Techniques: Similar optimization problems also occurred in the application of neural networks [79] that learn complex tasks in the field of AI4Science, e.g., approximating complex fields [80] and training on noisy measurements [81]. We believe the relevant research will benefit dynamic PAE generation and vice versa. In the field of reinforcement learning, research has been conducted on the problem of exploring the action space and utilizing inaccurate feedback [82]. However, the general dynamic PAE generation problem is defined in the complex state space, such as a patch with 10K+ pixels, which is intractable for reinforcement models. Recently, language-driven models have been adapted for optimization [83], but there is still a gap between language and low-level AE data [84], in which generative embedding models shall fill.
6 Discussions
Applications of the Framework: Although we only constructed a prototype system for physical attacks and test cases in a simulated environment during the experiment, our framework is general-purpose. Its applications include visual robustness evaluation of autonomous systems and building camouflage systems for interference or protecting key targets. Specifically, when evaluating the visual robustness, DynamicPAE provides guidance and efficiently improves the coverage of the physical-world-aligned adversarial testing. By incorporating existing technologies, such as dynamic materials (e.g., color-changing costumes [85]) and real-time 3D holography [86], DynamicPAE facilitates the construction of intelligent camouflage systems.
Limitations and Future Work: Our work focuses on the key generating algorithm for PAEs. However, further exploration is necessary to investigate additional properties, including the transferability of PAE between target models, the robustness of PAE on flexible materials, and the efficiency of the generator on more resource-limited devices. To achieve these properties, we can integrate existing black-box attacks, 3D simulations, and model light-weighting methods with the generative training method provided by DynamicPAE. While the efficacy of DynamicPAE in challenging learning tasks and its adoption of end-to-end modular design suggest the potential for effective integration, the characteristics of specific downstream tasks are still worth investigation. For the generative training technology, given the complexity of the PAE generator training task, more accurate models are needed to characterize the hardness of the exploration problem. The incorporation of technologies from related fields, as previously discussed, holds the potential for further enhancing PAE exploration.
7 Conclusion
@@
Future research directions can focus on technical improvements to the patch representation and training mechanism, exploring the application of the proposed framework, and facilitating the developed techniques in broader domains. In terms of application, although we have presented the preliminary prototype in our experiments, work on generating adversarial test and training data and exploring real-world red team attacks for specific tasks is still worthy of further investigation. We believe that our study has the broader potential to benefit the optimization of defense-insensitive adversarial attacks and noisy open-world tasks, and the model for the hardness of PAE optimization can bring insights to the defense of PAEs.
@@ Acknowledgments
This work is supported in part by Zhongguancun Laboratory, in part by State Key Laboratory of Complex & Critical Software Environment (CCSE), and in part by the National Natural Science Foundation of China under Grant 62506347.